OpenStudy (anonymous):

f(x)=2x - 5 ------- 4x + 2 g(x) = 2x^2 + 7 Find: f(g(x)) Please helpppp !!!!

OpenStudy (jkristia):

if f(x) = 2x - 5 and g(x) = 2x^2 + 7 and you have to find f(g(x)), then replace x in f(x) with 2x^2 + 7

OpenStudy (anonymous):

i dont get it?

OpenStudy (jkristia):

Not sure if I can explain this very well but... if the function is defined as f(x) = x^2 then f(z) = z^2 or f(q) = q^2 Meaning whichever is defined in the parentheses in the function 'f' is the variable given in the function itself, in this case <something>^2 so if f is given as f(1+x) then f(1+x) = (1+x)^2 since the function is <something>^2 so in case of f(g(x)), this means that any 'x' in the function defined by f(x) needs to be replaced by the definition of g(x) One thing I don't understand from your question is this ------- 4x + 2 What does this mean?